# On the Adversarial Risk of Test Time Adaptation: An Investigation into Realistic Test-Time Data Poisoning

### v2

@@ Abstract
Test-time adaptation (TTA) updates the model weights during the inference stage using testing data to enhance generalization. However, this practice exposes TTA to adversarial risks. Existing studies have shown that when TTA is updated with crafted adversarial test samples, also known as test-time poisoned data, the performance on benign samples can deteriorate. Nonetheless, the perceived adversarial risk may be overstated if the poisoned data is generated under overly strong assumptions. In this work, we first review realistic assumptions for test-time data poisoning, including white-box versus grey-box attacks, access to benign data, attack budget, and more. We then propose an effective and realistic attack method that better produces poisoned samples without access to benign samples, and derive an effective in-distribution attack objective. We also design two TTA-aware attack objectives. Our benchmarks of existing attack methods reveal that the TTA methods are more robust than previously believed. In addition, we analyze effective defense strategies to help develop adversarially robust TTA methods.
@@ 1 Introduction
Test-time adaptation (TTA) emerges as an effective measure to counter distribution shift at inference stage (Wang et al., 2020; Liu et al., 2021; Su et al., 2022; Song et al., 2023). Successful TTA methods leverage the testing data samples for self-training (Wang et al., 2020; Su et al., 2024b), distribution alignment Su et al. (2022); Liu et al. (2021) or prompt tuning (Gao et al., 2022). Despite the continuing efforts into developing computation efficient and high caliber TTA approaches, the robustness of TTA methods has not picked up until recently, leading to studies examining the robustness of TTA methods under constant distribution shift (Song et al., 2023), correlated testing data stream (Su et al., 2024a; Niu et al., 2023), open-world testing data (Li et al., 2023), adversarial robustness (Wu et al., 2023; Cong et al., 2023), etc. Among these risks, adversarial vulnerability warrants particular attention due to its potential for evading human inspection and the significant consequences of admitting malicious samples during TTA.
@@
Existing research frames the adversarial risk of Test-Time Adaptation (TTA) as the crafting of poisoned testing data, resulting in models updated with such data performing poorly on clean testing samples (Wu et al., 2023; Cong et al., 2023). Consequently, this task is also referred to as Test-Time Data Poisoning (TTDP). The pioneering work DIA (Wu et al., 2023) introduced a poisoning approach by crafting malicious data with access to all benign samples within a minibatch, leveraging real-time model weights for explicit gradient computing, i.e., a white-box attack. Another concurrent study (Cong et al., 2023) implements poisoning by preemptively injecting all poisoned data to attack the model even before TTA starts. While these explorations conclude that TTA methods are susceptible to poisoned data, evaluations based on unrealistic assumptions may exaggerate the adversarial risk for several reasons. i) Access to real-time model weights (white-box attack) is often considered overly optimistic, especially given that models are constantly updated during adaptation. Therefore, a grey-box or even black-box attack is preferred for TTDP. ii) The adversary is typically assumed only to be aware of the query samples submitted by themselves. Thus, benign samples submitted by other users should not be utilized for crafting poisoned data, for instance, through bi-level optimization (Wu et al., 2023). iii) Crafting poisoned data requires querying the model with testing samples. Repeatedly querying the model from a single user could easily trigger alerts in defensive systems. Therefore, any query sample, whether adversarial or benign, should be counted towards the attack budget. iv) Following the above concern, the adversary should not monopolize the entire testing bandwidth. This constraint translates to a scenario where poisoned data only partially occupies a minibatch, and the adversary is not allowed to inject all poisoned data at once even before TTA starts (Cong et al., 2023).
@@
  • We argue the unrealistic assumptions, e.g. white-box attack and overusing attack budget, made in existing attempts at TTDP may overestimate the adversarial risk of TTA methods. To address this, we first propose key criteria for defining realistic test-time data poisoning scenarios.
  • Under our proposed realistic test-time protocol, we analyze the generalization error and introduce an in-distribution attack strategy with feature distribution consistency regularization. This strategy eliminates the need for additional benign samples in evaluating the outer objective. Additionally, we tailor attack objectives specifically for TTA methods, resulting in more effective poisoning.
  • We conduct extensive evaluations on state-of-the-art TTA methods, demonstrating the efficacy of our proposed in-distribution attack strategy. Furthermore, we identify certain practices that are conducive to improving realistic adversarial robustness.
Figure 1: Illustration of the proposed test-time data poisoning pipeline. Adversary generates poisoned data by attacking a regularized objective without accessing benign samples from other users. Model is attacked when carrying out TTA on testing data stream mixed with benign and poisoned data.
@@ 2.1 Overview of Test-Time Data Poisoning
We first provide a generic overview of test-time adaptation for a $K$-way classification task. We denote the testing data as $\mathcal{D}=\{x_{i}\}_{i=1}^{N_{t}}$, a pre-trained model as $\theta_{0}$. TTA methods often employ unsupervised loss, $\mathcal{L}_{tta}$, to update model parameters upon observing a minibatch of testing samples $\mathcal{B}_{t}=\{x_{i}\}_{i=1}^{N_{b}}$ at timestamp $t$. Usually, a subset of model parameters $\theta_{t}^{u}\subseteq\theta_{t}$ is subject to update at timestamp $t$, and we denote the BN statistics as $\theta^{b}(\mathcal{B}_{t})=\{\mu(\mathcal{B}_{t}),\sigma^{2}(\mathcal{B}_{t})\}$ and the frozen parameters as $\theta^{f}$. The posterior of the sample $x_{i}$ towards the model parameter $\theta$ is $h(x_{i};\theta)\in[0,1]^{K}$. The adversarial risk arises when a subset of testing samples are poisoned, e.g. through adding an adversarial noise $\tilde{x}=x+\epsilon,\;s.t.\;||\epsilon||_{\infty}\leq b$. The model trained on poisoned data exhibit poor performance on clean/benign testing samples. In a typical TTA scenario, the online model is queried by both adversary and benign users. Thus, we denote the query data from adversary as adversary poisoned subset $\mathcal{B}_{a}=\{\tilde{x}_{i}\}$. The poisoned subset could be generated from arbitrary clean testing data, denoted as $\mathcal{B}_{ab}=\{x_{i}\}$ (e.g. use any public clean images). The generation follows an additive noise, i.e. $\tilde{x}_{i}=x_{i}+\epsilon_{i},\;s.t.\;\tilde{x}_{i}\in\mathcal{B}_{a}\;x_{i}\in\mathcal{B}_{ab}$, where the noise $\epsilon_{i}$ is the data poisoning to be learned. The query data from benign users is denoted as benign subset $\mathcal{B}_{b}=\{x_{i}\}$, where $\mathcal{B}_{ab}\cap\mathcal{B}_{b}=\emptyset$. The combination of both, $\mathcal{B}_{t}=\mathcal{B}_{a}\cup\mathcal{B}_{b}$, forms a single TTA minibatch. The effectiveness of test-time poisoning is evaluated at the attack success rate (classification error) on benign subset $\mathcal{B}_{b}$.
@@ 2.2 Realistic Test-Time Data Poisoning Protocol
White-Box v.s. Grey-Box Attack: Access to real-time TTA model parameters is a key factor for crafting realistic test-time data poisoning. Contrary to the assumption adopted by white-box adversarial attack (Szegedy et al., 2014; Biggio et al., 2013) that a frozen model is deployed for inference, the TTA model parameters experience constant updating at inference stage. The update is carried out on the cloud side, thus the attacker is assumed to have no access to real-time model weights. Such a realistic assumption prompts us to explore a more relaxed grey-box test-time data poisoning, i.e. only the model architecture and initial model weights are available to the adversary, such as the open-source famous pre-trained models (He et al., 2016; Dosovitskiy et al., 2021) and popular foundation models (Radford et al., 2021; Kirillov et al., 2023; Oquab et al., 2024).
Access To Benign Subset: The effectiveness of test-time data poisoning is evaluated on the benign subset $\mathcal{B}_{b}$ submitted by benign users. In a standard cloud service, users are generally restricted from accessing the queries of other users. Thus the adversary should only have access to adversary subset $\mathcal{B}_{a}$. This assumption prohibits the practice of crafting poisoned data by directly optimizing (minimizing) the loss on benign subset (Wu et al., 2023), on which the attack success rate (performance) is calculated.
@@
Setting Grey-box v.s. White-box Access Benign Subset Attack Order
@@
Adversarial Attacks on Standard Image Classification: Common grey-box or black-box adversarial attack techniques are often impractical for RTTDP due to two primary challenges. First, existing adversarial attack methods operate on a static model and inherently require multiple queries for gradient approximation or fitness evaluation, as seen in query-based attacks (Li et al., 2020; Xu et al., 2021), genetic algorithms (Chen et al., 2019), and black-box optimization (Qiu et al., 2021). However, in the test-time adaptation setting, the online model is continuously updated with each query during the inference phase. Thus, repetitive querying the model for gradient approximation or fitness evaluation is prohibited. Second, traditional adversarial attack methods focus on crafting adversarial samples to degrade the performance of the attacker’s own input data. In contrast, in a realistic test-time data poisoning scenario, poisoned samples are introduced into the test-time adaptation process to degrade the performance of benign samples submitted by other users.
@@ 3.1 Grey-Box Attack by Surrogate Model Distillation
To tackle the challenge of access to TTA model parameters, we propose to maintain a surrogate model, denoted as $\hat{\theta}_{t}$ at timestamp $t$, for the purpose of synthesizing poisoned data. To ensure good approximation, we distill the target model $\theta_{t}$ into the surrogate model $\hat{\theta}_{t}$ by leveraging the feedback of poisoned data from the online target model. Specifically, for each query to the target model, we minimize the symmetric KL-Divergence between the posteriors of the target and surrogate models, as Eq. 1. Our empirical observations demonstrated that utilizing the adversarial subset $\mathcal{B}_{a}$ for distillation yields performance comparable to a white-box attack, as illustrated in Fig. 2 (a).
@@ 3.2 In-Distribution Test-Time Data Poisoning
The above bi-level optimization problem employed in DIA (Wu et al., 2023) aims to generate adversarially poisoned data $\mathcal{B}_{a}$ by minimizing the loss function $\mathcal{L}_{atk}$, which is computed on the benign samples $\mathcal{B}_{b}$. Although DIA approximates $\theta_{t}^{u*}\approx\theta_{t}^{u}$ to discard the inner TTA gradient update loop and reduce the number of queries to the online model, several realistic concerns still persist under the RTTDP protocol. First, as discussed in Sec. 2.2, DIA, as a white-box attack method, must query the online TTA model $\theta^{*}_{t}$ for generating poisoned samples. To mitigate this issue, we propose to leverage a surrogate model $\hat{\theta}_{t}$ as an proxy model, which is distilled by Eq. 1 with the last feedback of $\mathcal{B}_{a,t-1}$. Second, DIA evaluates the outer optimization by employing the benign samples $\mathcal{B}_{b}$ (assuming access to benign users’ query samples). The adversarial risk mainly arises from the injection of poisoned samples into the TTA training process. However, employing the validation (benign) samples as the outer optimization objective may result in an overestimation of this risk, as the specific poisoned samples could be tailored for the attack on the inference of the current batch of benign samples (Park et al., 2024), e.g., ${\theta^{b}}^{\prime}$, rather than for attacking the TTA process, i.e., $\theta^{u}$.
@@
To tackle this challenge, we would like to approach from introducing additional constraints and combining the optimized and optimizing objectives into one objective to the problem, as Eq. 4, where $D(P_{1},P_{2})$ is a metric of two distributions and $P_{a}$ and $P_{ab}$ refer to the feature distributions of $\mathcal{B}_{a}$ and $\mathcal{B}_{ab}$. We have the follow reason why the formulation is effective. If $\mathcal{B}_{a}$ and $\mathcal{B}_{ab}$ have the similar distribution, we can have $P_{a}\rightarrow P_{ab}\Rightarrow\mathbbm{E}_{P_{a}(x)}[\mathcal{L}_{atk}(x)]\rightarrow\mathbbm{E}_{P_{ab}(x)}[\mathcal{L}_{atk}(x)]$ according to the Probably Approximately Correct (PAC) learning framework (Valiant, ). Thus attack against $\mathcal{B}_{a}$ has a high chance to generalize to $\mathcal{B}_{ab}$.
@@
With the introduced constraint $\mathcal{L}_{reg}=0$, we finally formulate the problem as Eq. 6. The problem now degenerates to a single level optimization with constraints which can be easily converted into an unconstrained optimization problem via Lagrangian multiplier (Lag, 2008). The unconstrained problem can be solved in an iterative fashion with detailed algorithm presented in the Appendix.
$\begin{split}&\min_{\mathcal{B}_{a}}\frac{1}{|\mathcal{B}_{a}|}\sum_{x_{i}\in\mathcal{B}_{a}}\mathcal{L}_{atk}\left(x_{i};\hat{\theta}_{t}^{\prime}(\mathcal{B}_{a})\right)\\ &s.t.\quad\hat{\theta}^{b\prime}=\{\mu(\mathcal{B}_{a}),\sigma^{2}(\mathcal{B}_{a})\};\quad\hat{\theta}_{t}^{\prime}(\mathcal{B}_{a})=\hat{\theta}_{t}^{u}\cup{\hat{\theta}^{b\prime}}\cup\hat{\theta}^{f};\quad\mathcal{L}_{reg}=0\\ \Rightarrow&\min_{\mathcal{B}_{a}}\max_{\lambda}\frac{1}{|\mathcal{B}_{a}|}\sum_{x_{i}\in\mathcal{B}_{a}}\mathcal{L}_{atk}\left(x_{i};\hat{\theta}_{t}^{\prime}(\mathcal{B}_{a})\right)+\lambda\mathcal{L}_{reg}\\ \Rightarrow&\min_{\mathcal{B}_{a}}\max_{\{\lambda_{0}\cdots\lambda_{L-1}\}}\frac{1}{|\mathcal{B}_{a}|}\sum_{x_{i}\in\mathcal{B}_{a}}\mathcal{L}_{atk}\left(x_{i};\hat{\theta}_{t}^{\prime}(\mathcal{B}_{a})\right)+\frac{1}{L}\sum_{l}\lambda_{l}KLD(\mathcal{N}(\mu_{i}^{l},\Sigma_{i}^{l})||\mathcal{N}(\tilde{\mu}_{i}^{l},\tilde{\Sigma}_{i}^{l}))\\ \end{split}$ (6)
Through optimizing the above objective, now we could craft the effective poisoned data $\mathcal{B}_{a}$ that satisfies $\mathbb{E}_{x_{i}\in\mathcal{B}_{a}}\mathcal{L}_{atk}(x_{i};\hat{\theta}_{t})\approx\mathbb{E}_{x_{i}\in\mathcal{B}_{b}}\mathcal{L}_{atk}(x_{i};\hat{\theta}_{t})$, since the approximation equation $P_{a}\rightarrow P_{ab}\rightarrow P_{b}$ now holds. Fig. 2 (c) demonstrates the effectiveness of this in-distribution attack.

### v3

@@ Abstract
Test-time adaptation (TTA) updates the model weights during the inference stage using testing data to enhance generalization. However, this practice exposes TTA to adversarial risks. Existing studies have shown that when TTA is updated with crafted adversarial test samples, also known as test-time poisoned data, the performance on benign samples can deteriorate. Nonetheless, the perceived adversarial risk may be overstated if the poisoned data is generated under overly strong assumptions. In this work, we first review realistic assumptions for test-time data poisoning, including white-box versus grey-box attacks, access to benign data, attack order, and more. We then propose an effective and realistic attack method that better produces poisoned samples without access to benign samples, and derive an effective in-distribution attack objective. We also design two TTA-aware attack objectives. Our benchmarks of existing attack methods reveal that the TTA methods are more robust than previously believed. In addition, we analyze effective defense strategies to help develop adversarially robust TTA methods. The source code is available at https://github.com/Gorilla-Lab-SCUT/RTTDP.
**footnotetext: Equal contribution. Correspondence to <xu_xun@i2r.a-star.edu.sg>. This work was done during Yongyi Su, Yushu Li and Nanqing Liu’s visit to I2R.
@@ 1 Introduction
Test-time adaptation (TTA) emerges as an effective measure to counter distribution shift at inference stage (Wang et al., 2020; Su et al., 2022; Zhong et al., 2022; Chen et al., 2023; Chi et al., 2024; Wu et al., 2024). Successful TTA methods leverage the testing data samples for self-training (Wang et al., 2020; Su et al., 2024b), distribution alignment (Liu et al., 2021; Su et al., 2022) or prompt tuning (Gao et al., 2022; Liu et al., 2025). Despite the continuing efforts into developing computation efficient and high caliber TTA approaches, the robustness of TTA methods has not picked up until recently, leading to studies examining the robustness of TTA methods under constant distribution shift (Song et al., 2023), correlated testing data stream (Su et al., 2024a; Niu et al., 2023), open-world testing data (Li et al., 2023), adversarial robustness (Wu et al., 2023; Cong et al., 2023), etc. Among these risks, adversarial vulnerability warrants particular attention due to its potential for evading human inspection and the significant consequences of admitting malicious samples during TTA.
@@
Existing research frames the adversarial risk of Test-Time Adaptation (TTA) as the crafting of poisoned testing data, resulting in models updated with such data performing poorly on clean testing samples (Wu et al., 2023; Cong et al., 2023). Consequently, this task is also referred to as Test-Time Data Poisoning (TTDP). The pioneering work DIA (Wu et al., 2023) introduced a poisoning approach by crafting malicious data with access to all benign samples within a minibatch, leveraging real-time model weights for explicit gradient computing, i.e., a white-box attack. Another concurrent study (Cong et al., 2023) implements poisoning by preemptively injecting all poisoned data to attack the model even before TTA starts. While these explorations conclude that TTA methods are susceptible to poisoned data, evaluations based on unrealistic assumptions may exaggerate the adversarial risk for several reasons. i) Access to real-time model weights (white-box attack) is often considered overly optimistic, especially given that models are constantly updated during adaptation. Therefore, a grey-box or even black-box attack is preferred for TTDP. ii) The adversary is typically assumed only to be aware of the query samples submitted by themselves. Thus, benign samples submitted by other users should not be utilized for crafting poisoned data, for instance, through bi-level optimization (Wu et al., 2023). iii) Crafting poisoned data requires querying the model with testing samples (Wu et al., 2023). Repeatedly querying the model from a single user could easily trigger alerts in defensive systems. Therefore, any query sample, whether adversarial or benign, should be counted towards the attack budget. iv) Following the above concern, the adversary should not monopolize the entire testing bandwidth. This constraint translates to a scenario where poisoned data only partially occupies the testing stream, and the adversary is not allowed to inject all poisoned data at once even before TTA starts (Cong et al., 2023).
@@
  • We argue the unrealistic assumptions, e.g. white-box attack, access to benign subset and offline attack order, made in existing attempts at TTDP may overestimate the adversarial risk of TTA methods. To address this, we first propose key criteria for defining realistic test-time data poisoning scenarios.
  • Under our proposed realistic test-time protocol, we analyze the generalization error and introduce an in-distribution attack strategy with feature consistency regularization. This strategy eliminates the need for additional benign samples in evaluating the outer objective. In addition, we tailor attack objectives specifically for TTA methods, resulting in more effective poisoning.
  • We perform extensive evaluations on state-of-the-art TTA methods, demonstrating the efficacy of our proposed in-distribution attack strategy. Furthermore, we identify certain practices that are conducive to improving realistic adversarial robustness.
Figure 1: Illustration of the proposed Realistic Test-Time Data Poisoning (RTTDP) pipeline. $\mathcal{B}_{ab}$ indicates the adversary benign subset, and $\mathcal{B}_{a}$ indicates the adversary poisoned subset where the samples are poisoned from the clean samples in $\mathcal{B}_{ab}$. $\mathcal{B}_{b}$ indicates the benign users’ subset where the samples are used to validate the adversarial risk of TTA pipeline and these samples cannot be access by the adversary. Adversary generates poisoned data by attacking a regularized objective without accessing benign samples from other users. Model is attacked when carrying out TTA on testing data stream mixed with benign and poisoned data.
2 Related Work
Test-Time Adaptation (TTA): Wang et al. (2020; 2022); Niu et al. (2022); Su et al. (2022); Li et al. (2023); Liang et al. (2024); Wu et al. (2024) and Wang et al. (2025) have shown significant success in bridging the domain gap by using stream-based testing samples to dynamically update models in real time. The success of TTA is mainly attributed to the self-supervised learning on testing data. While it has proved sensitive to confirmation bias (Arazo et al., 2020), many solutions were proposed to minimize the influences of the wrong pseudo-labels, including minimizing sample entropy (Wang et al., 2020; Liang et al., 2020), adding regularization terms (Song et al., 2023; Su et al., 2024b), using confidence thresholding (Niu et al., 2022; 2023), updating models with exponential moving average architectures (Wang et al., 2022; Döbler et al., 2023), partially updating model weights (Wang et al., 2020; Yuan et al., 2023), and augmenting testing samples (Zhang et al., 2022; Döbler et al., 2023), all aimed at minimizing sample distribution discrepancies. However, adaptation during the testing stage remains highly risky. In this work, we aim to investigate the risk of TTA posed by data poisoning.
Adversarial Attack & Data Poisoning: Adversarial risk is a crucial concern for models to address for safe deployment, which can be divided into two categories: adversarial attacks and data poisoning. Adversarial attacks (Szegedy et al., 2014; Akhtar & Mian, 2018; Goodfellow et al., 2014; Madry et al., 2018; Croce & Hein, 2020; Chen et al., 2022; Chakraborty et al., 2018) manipulate models during inference by adding small perturbations to input data. White-box attacks (Szegedy et al., 2014; Tramèr et al., 2018) assume full access to the victim model, creating adversarial examples by maximizing loss gradients. In contrast, grey-box attacks (Chen et al., 2017; Ilyas et al., 2018; Ru et al., 2019) assume no model access, generating adversarial samples by estimating gradients through intensive querying. Data poisoning (Biggio et al., 2012; Yang et al., 2017; Shafahi et al., 2018; Alfeld et al., 2016; Huang et al., 2021; Fowl et al., 2021; Fan et al., 2022) compromises models by injecting manipulated data into the training set, misleading the training process. Traditional data poisoning assumes that the attacker can observe and poison the entire training set at once. Recently, online poisoning (Zhang et al., 2020) relaxes this assumption by requiring knowledge of the model updating strategies. In this work, we explore a more realistic scenario of TTA’s adversarial robustness, focusing on data poisoning without online model access and without multiple queries for the same data.
@@ 3.1 Overview of Test-Time Data Poisoning
We first provide a generic overview of test-time adaptation for a $K$-way classification task. We denote the testing data as $\mathcal{D}=\{x_{i}\}_{i=1}^{N_{t}}$ and a pre-trained model as $\theta_{0}$. TTA methods often employ unsupervised loss, $\mathcal{L}_{tta}$, to update model parameters upon observing a minibatch of testing samples $\mathcal{B}_{t}=\{x_{i}\}_{i=1}^{N_{b}}$ at timestamp $t$. Usually, a subset of model parameters $\theta_{t}^{u}\subseteq\theta_{t}$ is subject to update at timestamp $t$, and we denote the BN statistics as $\theta^{b}(\mathcal{B}_{t})=\{\mu(\mathcal{B}_{t}),\sigma^{2}(\mathcal{B}_{t})\}$ and the frozen parameters as $\theta^{f}$. The posterior of the sample $x_{i}$ towards the model parameter $\theta$ is $h(x_{i};\theta)\in[0,1]^{K}$. The adversarial risk arises when a subset of testing samples are poisoned, e.g. through adding an adversarial noise $\tilde{x}=x+\epsilon,\;s.t.\;||\epsilon||_{\infty}\leq b$. The model trained on poisoned data exhibit poor performance on clean/benign testing samples. In a typical TTA scenario, the online model is queried by both adversary and benign users. Thus, we denote the query data from adversary as adversary poisoned subset $\mathcal{B}_{a}=\{\tilde{x}_{i}\}$. The poisoned subset could be generated from arbitrary clean testing data, denoted as $\mathcal{B}_{ab}=\{x_{i}\}$ (e.g. use any public clean images). The generation follows an additive noise, i.e. $\tilde{x}_{i}=x_{i}+\epsilon_{i},\;s.t.\;\tilde{x}_{i}\in\mathcal{B}_{a}\;x_{i}\in\mathcal{B}_{ab}$, where the noise $\epsilon_{i}$ is the data poisoning to be learned. The query data from benign users is denoted as benign subset $\mathcal{B}_{b}=\{x_{i}\}$, where $\mathcal{B}_{ab}\cap\mathcal{B}_{b}=\emptyset$. Generally, we form the combination of both, $\mathcal{B}_{t}=\mathcal{B}_{a}\cup\mathcal{B}_{b}$, as a single TTA minibatch. The effectiveness of test-time poisoning is evaluated at the attack success rate (classification error) on benign subset $\mathcal{B}_{b}$. An illustration of batch split is presented in the Appendix A.1.1.
@@ 3.2 Realistic Test-Time Data Poisoning Protocol
White-Box v.s. Grey-Box Attack: Access to real-time TTA model parameters is a key factor for crafting realistic test-time data poisoning. Contrary to the assumption adopted by white-box adversarial attack (Szegedy et al., 2014; Biggio et al., 2013) that a frozen model is deployed for inference, the TTA model parameters experience constant updating at inference stage. The update is performed on the cloud side, thus the attacker doesn’t normally have access to real-time model weights. Such a realistic assumption prompts us to explore a more relaxed grey-box test-time data poisoning, i.e. only the model architecture and initial model weights are available to the adversary, such as the open-source famous pre-trained models (He et al., 2016; Dosovitskiy et al., 2021) and popular foundation models (Radford et al., 2021; Kirillov et al., 2023; Oquab et al., 2024).
Access To Benign Subset: The effectiveness of TTDP is evaluated on the benign subset $\mathcal{B}_{b}$ submitted by benign users. In a standard cloud service, users are generally restricted from accessing the queries of other users. Thus the adversary should only have access to adversary subset $\mathcal{B}_{a}$. This assumption prohibits the practice of crafting poisoned data by directly optimizing (minimizing) the loss on benign subset (Wu et al., 2023), on which the attack success rate (performance) is calculated.
@@
Setting Grey-box v.s. White-box Access to Benign Subset Attack Order
@@
Adversarial Attacks on Standard Image Classification: Common grey-box or black-box adversarial attack techniques are often impractical for RTTDP due to two primary challenges. First, existing adversarial attack methods operate on a static model and inherently require multiple queries for gradient approximation or fitness evaluation, as seen in query-based attacks (Li et al., 2020; Xu et al., 2021), genetic algorithms (Chen et al., 2019), and black-box optimization (Qiu et al., 2021). However, in the test-time adaptation setting, the online model is continuously updated with each query during the inference phase. Thus, repetitive querying the model for gradient approximation or fitness evaluation is unavailable. Second, traditional adversarial attack methods focus on crafting adversarial samples to degrade the performance of the attacker’s own input data. In contrast, in a realistic test-time data poisoning scenario, poisoned samples are introduced into the test-time adaptation process to degrade the performance of benign samples submitted by other users.
@@ 4.1 Grey-Box Attack by Surrogate Model Distillation
To tackle the challenge of access to TTA model parameters, we propose to maintain a surrogate model, denoted as $\hat{\theta}_{t}$ at timestamp $t$, for the purpose of synthesizing poisoned data. To ensure good approximation, we distill the target model $\theta_{t}$ into the surrogate model $\hat{\theta}_{t}$ by leveraging the feedback of poisoned data from the online target model. Specifically, for each query to the target model, we minimize the symmetric KL-Divergence between the posteriors of the target and surrogate models, as Eq. 1. Our empirical observations demonstrate that utilizing the adversarial subset $\mathcal{B}_{a}$ for distillation yields performance comparable to a white-box attack, as illustrated in Fig. 2 (a).
@@ 4.2 In-Distribution Test-Time Data Poisoning
The above bi-level optimization problem employed in DIA (Wu et al., 2023) aims to generate adversarially poisoned data $\mathcal{B}_{a}$ by minimizing the loss function $\mathcal{L}_{atk}$, which is computed on the benign samples $\mathcal{B}_{b}$. Although DIA approximates $\theta_{t}^{u*}\approx\theta_{t}^{u}$ to discard the inner TTA gradient update loop and reduce the number of queries to the online model, several realistic concerns still persist under the RTTDP protocol. First, as discussed in Sec. 3.2, DIA, as a white-box attack method, must query the online TTA model $\theta^{*}_{t}$ for generating poisoned samples. To mitigate this issue, we propose to leverage a surrogate model $\hat{\theta}_{t}$ as an proxy model, which is distilled by Eq. 1 with the last feedback of $\mathcal{B}_{a,t-\delta}$, where $\delta$ denotes the time interval between two injected poisoned subsets. Second, DIA evaluates the outer optimization by employing the benign samples $\mathcal{B}_{b}$ (assuming access to benign users’ query samples). The adversarial risk mainly arises from the injection of poisoned samples into the TTA training process. However, employing the validation (benign) samples as the outer optimization objective may result in an overestimation of this risk, as the specific poisoned samples could be tailored for the attack on the inference of the current batch of benign samples (Park et al., 2024), e.g., ${\theta^{b\prime}}$, rather than for attacking the TTA process, i.e., $\theta^{u}$.
@@
To address this challenge, we propose introducing additional constraints and integrating the optimized target $\mathcal{B}_{ab}$ with the optimizing objective $\mathcal{B}_{a}$ into a unified objective $\mathcal{B}_{a}$, as presented in Eq. 4, where $D(P_{1},P_{2})$ is a metric of two distributions and $P_{a}$ and $P_{ab}$ refer to the feature distributions of $\mathcal{B}_{a}$ and $\mathcal{B}_{ab}$. We have the follow reason why the formulation is effective. If $\mathcal{B}_{a}$ and $\mathcal{B}_{ab}$ have the similar distribution, we can have $P_{a}\rightarrow P_{ab}\Rightarrow\mathbbm{E}_{P_{a}(x)}[\mathcal{L}_{atk}(x)]\rightarrow\mathbbm{E}_{P_{ab}(x)}[\mathcal{L}_{atk}(x)]$ according to the Probably Approximately Correct (PAC) learning framework (Valiant, ). Thus attack against $\mathcal{B}_{a}$ has a high chance to generalize to $\mathcal{B}_{ab}$.
@@
With the introduced constraint $\mathcal{L}_{reg}=0$, we finally formulate the problem as Eq. 6. The problem now degenerates to a single level optimization with constraints which can be easily converted into an unconstrained optimization problem via Lagrangian multiplier (Lag, 2008). The unconstrained problem can be solved in an iterative fashion with detailed algorithm presented in the Appendix A.1.6.
$\begin{split}&\min_{\mathcal{B}_{a}}\frac{1}{|\mathcal{B}_{a}|}\sum_{x_{i}\in\mathcal{B}_{a}}\mathcal{L}_{atk}\left(x_{i};\hat{\theta}_{t}^{\prime}(\mathcal{B}_{a})\right)\\ &s.t.\quad\hat{\theta}^{b\prime}=\{\mu(\mathcal{B}_{a}),\sigma^{2}(\mathcal{B}_{a})\};\quad\hat{\theta}_{t}^{\prime}(\mathcal{B}_{a})=\hat{\theta}_{t}^{u}\cup{\hat{\theta}^{b\prime}}\cup\hat{\theta}^{f};\quad\mathcal{L}_{reg}=0\\ \Rightarrow&\min_{\mathcal{B}_{a}}\max_{\lambda}\frac{1}{|\mathcal{B}_{a}|}\sum_{x_{i}\in\mathcal{B}_{a}}\mathcal{L}_{atk}\left(x_{i};\hat{\theta}_{t}^{\prime}(\mathcal{B}_{a})\right)+\lambda\mathcal{L}_{reg}\\ \Rightarrow&\min_{\mathcal{B}_{a}}\max_{\{\lambda_{0}\cdots\lambda_{L-1}\}}\frac{1}{|\mathcal{B}_{a}|}\sum_{x_{i}\in\mathcal{B}_{a}}\mathcal{L}_{atk}\left(x_{i};\hat{\theta}_{t}^{\prime}(\mathcal{B}_{a})\right)+\frac{1}{L}\sum_{l}\lambda_{l}KLD(\mathcal{N}(\mu_{i}^{l},\Sigma_{i}^{l})||\mathcal{N}(\tilde{\mu}_{i}^{l},\tilde{\Sigma}_{i}^{l})),\\ \end{split}$ (6)
where $L$ is the number of intermediate feature layers. Through optimizing the above objective, now we could craft the effective poisoned data $\mathcal{B}_{a}$ that satisfies $\mathbb{E}_{x_{i}\in\mathcal{B}_{a}}\mathcal{L}_{atk}(x_{i};\hat{\theta}_{t})\approx\mathbb{E}_{x_{i}\in\mathcal{B}_{b}}\mathcal{L}_{atk}(x_{i};\hat{\theta}_{t})$, since the approximation equation $P_{a}\rightarrow P_{ab}\rightarrow P_{b}$ now holds. Fig. 2 (c) demonstrates the effectiveness of this in-distribution attack, where the attack performance of the attack objective combined with the feature regularization (orange bars) is always higher than that of the corresponding attack objective alone (blue bars) in different TTA methods. Next, we would introduce the attack objectives $\mathcal{L}_{atk}$ that we designed to effectively generate different kinds of poisoned samples.